# Trial and Trust: Addressing Byzantine Attacks with Comprehensive Defense Strategy

### v1

@@ F.1  Scaled BANT
10:     $\omega^{t}=(1-\beta)\omega_{i}^{t-1}+\beta\frac{[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}g_{i}^{t})]_{0}}{\sum\limits_{j=1}^{n}[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}g_{j}^{t})]_{0}}$
@@
14:     $x^{t+1}=x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\limits_{i=1}^{n}\mathbb{I}_{[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}g_{i}^{t})>0]}\omega_{i}^{t}g_{i}^{t}$
@@
Under Assumptions 2.1, 2.2(b), 2.5, F.1, the following holds for the iteration of Algorithm 5:
@@
$-\gamma\left\langle\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}$ $=$ $\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}-\gamma\left\langle\nabla f(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}.$
@@
$\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}$ (${\leqslant}$ $\frac{\gamma}{2}\left\|\nabla f(x)-\nabla\hat{f}(x)\right\|_{\hat{P}^{-1}}^{2}+\frac{\gamma}{2}\left\|\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|_{\hat{P}^{-1}}^{2}$
@@
$-\gamma\left\langle\nabla f(x),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}$ ${=}$ $-\frac{\gamma}{2}\|\nabla f(x)\|_{\hat{P}^{-1}}^{2}-\frac{\gamma}{2}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|_{\hat{P}^{-1}}^{2}$
@@
(${\leqslant}$ $-\frac{\gamma}{2}\|\nabla f(x)\|_{\hat{P}^{-1}}^{2}-\frac{\gamma}{2}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x)\right\|_{\hat{P}^{-1}}^{2}$
@@
Under Assumptions 2.1, 2.2(b), 2.3, 2.4 with $\delta_{2}\leqslant\frac{2\Gamma-\alpha}{\alpha+\frac{4\Gamma^{2}}{\alpha^{2}}}$, 2.5, F.1, for solving the problem (1), after $T$ iteration of Algorithm 5 with $\gamma\leqslant\frac{\alpha}{12L}$, it holds that
@@
$\mathbb{E}\hat{f}(x^{t+1})$$\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\gamma\beta\cdot\frac{G}{n}\left\langle\nabla\hat{f}(x^{t}),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x^{t})\right\rangle_{(\hat{P}^{t})^{-1}}+\frac{L\gamma^{2}\beta}{2n\alpha^{2}}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}\right\|^{2}$
@@
(${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\frac{\gamma\beta G}{2n\alpha}\zeta(N)-\frac{\gamma\beta G}{2n\Gamma}\left\|\nabla f(x^{t})\right\|^{2}+\frac{\gamma\beta G}{2n\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}\beta}{2n\alpha^{2}}\left(\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f(x^{t})-f_{i}(x^{t})\right\|^{2}+\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x^{t})-g_{i}^{t}\right\|^{2}\right)+\frac{3L\gamma^{2}\beta G}{2n\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}$
@@
We get a result similar to Corollary 3.2, but with the deterioration that each term is multiplied by an additional constant $\nicefrac{{\Gamma}}{{\alpha}}>1$. This result suits us, since in (Sadiev et al., 2024) the result for the Scaled SARAH method corresponds similarly to the result for the classical SARAH (Nguyen et al., 2017) method.
@@ F.2  Scaled AutoBANT
10:     $\omega^{t}\approx\arg\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)$
11:     $x^{t+1}=x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$
@@
Under Assumptions 2.1, 2.2(c) 2.3, 2.4 with $\delta_{2}\leqslant 0.25$, 2.5, F.1, for solving the problem (1), after $T$ iterations of Algorithm 6 with $\gamma\leqslant\frac{\alpha}{12L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]\cdot 4\Gamma}{\gamma T}+\frac{3\Gamma}{\alpha}\delta_{1}+\frac{6L\gamma\Gamma}{\alpha^{2}G}\sigma^{2}+\frac{2\Gamma}{\alpha}\zeta(N)+4\Gamma\frac{\delta}{\gamma}.$
@@
$\hat{f}(x^{t+1})$$\leqslant$ $\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma(\hat{P}^{t})^{-1}\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)+\delta$
$\leqslant$ $\hat{f}\left(x^{t}-\gamma(\hat{P}^{t})^{-1}\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right)+\delta$
(${\leqslant}$ $\hat{f}\left(x^{t}\right)-\gamma\left\langle\nabla\hat{f}(x^{t}),(\hat{P}^{t})^{-1}\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\rangle+\frac{L\gamma^{2}}{2}\left\|(\hat{P}^{t})^{-1}\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
@@
$\mathbb{E}\hat{f}(x^{t+1})$$\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\gamma\left\langle\nabla\hat{f}(x^{t}),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x^{t})\right\rangle_{(\hat{P}^{t})^{-1}}+\frac{L\gamma^{2}}{2\alpha^{2}}\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
@@
$+\frac{\gamma}{2\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}}{2\alpha^{2}}\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}+\delta$
(${\leqslant}$ $\mathbb{E}\hat{f}(x^{t})+\frac{\gamma}{2\alpha}\zeta(N)-\frac{\gamma}{2\Gamma}\left\|\nabla f(x^{t})\right\|^{2}+\frac{\gamma}{2\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2\alpha^{2}}\left(\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}(\nabla f(x^{t})-f_{i}(x^{t}))\right\|^{2}+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla(f_{i}(x^{t})-g_{i}^{t})\right\|^{2}\right)$
$+\frac{3L\gamma^{2}}{2\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}+\delta.$
@@
$+\frac{3L\gamma^{2}}{2G\alpha^{2}}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|(\nabla f(x^{t})-f_{i}(x^{t}))\right\|^{2}+\frac{3L\gamma^{2}}{2G^{2}\alpha^{2}}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla(f_{i}(x^{t})-g_{i}^{t})\right\|^{2}$
$+\frac{3L\gamma^{2}}{2\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
@@
$+\frac{3L\gamma^{2}}{2\alpha^{2}}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\frac{1}{G}\sigma^{2}\right)+\frac{3L\gamma^{2}}{2\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}+\delta$
@@
$+\frac{\gamma}{2\alpha}(1+\frac{3L\gamma}{\alpha})\delta_{1}+\frac{3L\gamma^{2}}{2\alpha^{2}G}\sigma^{2}+\frac{\gamma}{2\alpha}\zeta(N)+\delta.$
@@
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]\cdot 4\Gamma}{\gamma T}+\frac{3\Gamma}{\alpha}\delta_{1}+\frac{6L\gamma\Gamma}{\alpha^{2}G}\sigma^{2}+\frac{2\Gamma}{\alpha}\zeta(N)+4\Gamma\frac{\delta}{\gamma}.$
@@
$\frac{1}{T}\sum\limits_{i=1}^{T-1}\|\nabla f(x^{t})\|^{2}$ $=$ $\mathcal{O}\Biggl{(}\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]LG}{T}\cdot\frac{\Gamma}{\alpha}+\frac{\sigma\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]LG}}{\sqrt{T}}\cdot\frac{\Gamma}{\alpha}+\left(\delta_{1}+\zeta(N)\right)\cdot\frac{\Gamma}{\alpha}$
$+\delta\Biggl{(}L+\frac{\sqrt{TL}\sigma}{\sqrt{\mathbb{E}[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})]G}}\Biggr{)}\cdot\frac{\Gamma}{\alpha}\Biggr{)}$
@@ Appendix G Local methods
As highlighted in the main section, the significant expense associated with communication remains a critical concern in various fields. The communication bottleneck can act as a substantial barrier, limiting efficiency and hindering progress. To address this challenge, many researches are turning to local approaches, which focus on minimizing the need for exchanging the information (Woodworth et al., 2020; Koloskova et al., 2020; Khaled et al., 2020; Gorbunov et al., 2021b). The idea is that each device performs a predefined number of local steps without utilizing information from other devices, and at the end of such a round, the server performs a mutual update. In this section we adapt our AutoBANT algorithm to this scenario. Below we present the formal description of the Local AutoBANT method (Algorithm 7).
@@
Under Assumptions 2.1, 2.2(b), 2.3, 2.4, 2.5, at each iteration $t$ of Algorithm 7 with $\gamma\leqslant\frac{1}{4(l-1)L}$, the following estimate is valid:
@@
$\mathbb{E}\left\|x_{i}^{t+1}-\overline{x}^{t+1}\right\|^{2}$ $=$ $\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}+\gamma^{2}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-2\gamma\mathbb{E}\left\langle x_{i}^{t}-\overline{x}^{t},g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\rangle$
$=$ $\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}+\gamma^{2}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-2\gamma\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle.$
@@
$\mathbb{E}V^{t+1}$ $=$ $\mathbb{E}V^{t}+\frac{\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-\frac{2\gamma}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle+2\gamma\left\langle\overline{x}^{t}-\overline{x}^{t},\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle$ (7)
@@
=$\overset{(i)}{=}$ $\mathbb{E}\left\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right\|^{2}+\mathbb{E}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+2\mathbb{E}\left\langle g_{i}^{t}-\nabla f_{i}(x_{i}^{t}),\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
=$\overset{(ii)}{=}$ $\mathbb{E}\left\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right\|^{2}+\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$ $=$ $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right\|^{2}+\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$ (8)
$+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-2\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$\overset{(i)}{\leqslant}$ $\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}+\sigma^{2},$

### v5

@@ F.1 Scaled Bant
10:  $\omega^{t}=(1-\beta)\omega_{i}^{t-1}+\beta\frac{[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}g_{i}^{t})]_{0}}{\sum\nolimits_{j=1}^{n}[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}g_{j}^{t})]_{0}}$
@@
14:  $x^{t+1}=x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\nolimits_{i=1}^{n}\mathbb{I}_{[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}g_{i}^{t})>0]}\omega_{i}^{t}g_{i}^{t}$
@@
Under Assumptions 3.1, 3.2(b), 3.5, F.1, the following holds for the iteration of Algorithm 5:
@@
$-\gamma\left\langle\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}$ $=$ $\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}$
$-\gamma\left\langle\nabla f(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}.$
@@
$\gamma\left\langle\nabla f(x)-\nabla\hat{f}(x),\frac{1}{G}\sum_{i\in\mathcal{G}}\nabla f_{i}(x)\right\rangle_{\hat{P}^{-1}}$
[…]
@@
[…]
@@
[…]
@@
Under Assumptions 3.1, 3.2(b), 3.3, 3.4 with $\delta_{2}\leqslant\frac{2\Gamma-\alpha}{\alpha+\frac{4\Gamma^{2}}{\alpha^{2}}}$, 3.5, F.1, for solving the problem (1), after $T$ iteration of Algorithm 5 with $\gamma\leqslant\frac{\alpha}{12L}$, it holds that
@@
$\mathbb{E}\hat{f}(x^{t+1})$ $\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\gamma\beta\cdot\frac{G}{n}\left\langle\nabla\hat{f}(x^{t}),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x^{t})\right\rangle_{(\hat{P}^{t})^{-1}}$
$+\frac{L\gamma^{2}\beta}{2n\alpha^{2}}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}\right\|^{2}$
@@
[…]
$+\frac{3L\gamma^{2}\beta}{2n\alpha^{2}}\left(\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f(x^{t})-f_{i}(x^{t})\right\|^{2}+\sum_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla f_{i}(x^{t})-g_{i}^{t}\right\|^{2}\right)$
$+\frac{3L\gamma^{2}\beta G}{2n\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}$
@@
We get a result similar to Corollary 4.2, but with the deterioration that each term is multiplied by an additional constant $\nicefrac{{\Gamma}}{{\alpha}}>1$. This result suits us, since in [sadiev2024stochastic] the result for the Scaled SARAH method corresponds similarly to the result for the classical SARAH [nguyen2017sarah] method.
@@ F.2 Scaled AutoBant
10:  $\omega^{t}\approx\arg\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\nolimits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)$
11:  $x^{t+1}=x^{t}-\gamma\left(\hat{P}^{t}\right)^{-1}\sum\nolimits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$
@@
Under Assumptions 3.1, 3.2(c) 3.3, 3.4 with $\delta_{2}\leqslant 0.25$, 3.5, F.1, for solving the problem (1), after $T$ iterations of Algorithm 6 with $\gamma\leqslant\frac{\alpha}{12L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$ $\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]\cdot 4\Gamma}{\gamma T}+\frac{3\Gamma}{\alpha}\delta_{1}+\frac{6L\gamma\Gamma}{\alpha^{2}G}\sigma^{2}+\frac{2\Gamma}{\alpha}\zeta(N).$
@@
$\hat{f}(x^{t+1})$ $=$ $\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma(\hat{P}^{t})^{-1}\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)$
$\leqslant$ $\hat{f}\left(x^{t}-\gamma(\hat{P}^{t})^{-1}\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right)$
[…]
$+\frac{L\gamma^{2}}{2}\left\|(\hat{P}^{t})^{-1}\frac{1}{G}\sum_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$\mathbb{E}\hat{f}(x^{t+1})$ $\leqslant$ $\mathbb{E}\hat{f}(x^{t})-\gamma\left\langle\nabla\hat{f}(x^{t}),\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x^{t})\right\rangle_{(\hat{P}^{t})^{-1}}$
$+\frac{L\gamma^{2}}{2\alpha^{2}}\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$+\frac{\gamma}{2\alpha}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}\right)+\frac{L\gamma^{2}}{2\alpha^{2}}\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
[…]
$+\frac{3L\gamma^{2}}{2\alpha^{2}}\left(\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}(\nabla f(x^{t})-f_{i}(x^{t}))\right\|^{2}\right$.
$\left.+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla(f_{i}(x^{t})-g_{i}^{t})\right\|^{2}\right)+\frac{3L\gamma^{2}}{2\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}.$
@@
$+\frac{3L\gamma^{2}}{2G\alpha^{2}}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|(\nabla f(x^{t})-f_{i}(x^{t}))\right\|^{2}$
$+\frac{3L\gamma^{2}}{2G^{2}\alpha^{2}}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|\nabla(f_{i}(x^{t})-g_{i}^{t})\right\|^{2}+\frac{3L\gamma^{2}}{2\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}$
@@
$+\frac{3L\gamma^{2}}{2\alpha^{2}}\left(\delta_{1}+\delta_{2}\|\nabla f(x^{t})\|^{2}+\frac{1}{G}\sigma^{2}\right)+\frac{3L\gamma^{2}}{2\alpha^{2}}\left\|\nabla f(x^{t})\right\|^{2}$
@@
$+\frac{\gamma}{2\alpha}(1+\frac{3L\gamma}{\alpha})\delta_{1}+\frac{3L\gamma^{2}}{2\alpha^{2}G}\sigma^{2}+\frac{\gamma}{2\alpha}\zeta(N).$
@@
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$ $\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]\cdot 4\Gamma}{\gamma T}+\frac{3\Gamma}{\alpha}\delta_{1}+\frac{6L\gamma\Gamma}{\alpha^{2}G}\sigma^{2}+\frac{2\Gamma}{\alpha}\zeta(N).$
@@
$\frac{1}{T}\sum\limits_{i=1}^{T-1}\|\nabla f(x^{t})\|^{2}=$ $\mathcal{O}\Biggl(\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]LG}{T}\cdot\frac{\Gamma}{\alpha}+\frac{\sigma\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]LG}}{\sqrt{T}}\cdot\frac{\Gamma}{\alpha}+\left(\delta_{1}+\zeta(N)\right)\cdot\frac{\Gamma}{\alpha}\Biggr)$
@@ Appendix G Local methods
As highlighted in the main section, the significant expense associated with communication remains a critical concern in various fields. The communication bottleneck can act as a substantial barrier, limiting efficiency and hindering progress. To address this challenge, many researches are turning to local approaches, which focus on minimizing the need for exchanging the information [woodworth2020minibatch, koloskova2020unified, khaled2020tighter, gorbunov2021local]. The idea is that each device performs a predefined number of local steps without utilizing information from other devices, and at the end of such a round, the server performs a mutual update. In this section we adapt our AutoBant algorithm to this scenario. Below we present the formal description of the Local AutoBant method (Algorithm 7).
@@
Under Assumptions 3.1, 3.2(b), 3.3, 3.4, 3.5, at each iteration $t$ of Algorithm 7 with $\gamma\leqslant\frac{1}{4(l-1)L}$, the following estimate is valid:
@@
$\mathbb{E}\left\|x_{i}^{t+1}-\overline{x}^{t+1}\right\|^{2}$ $=$ $\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}+\gamma^{2}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
$-2\gamma\mathbb{E}\left\langle x_{i}^{t}-\overline{x}^{t},g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\rangle$
$=$ $\mathbb{E}\left\|x_{i}^{t}-\overline{x}^{t}\right\|^{2}+\gamma^{2}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
$-2\gamma\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle.$
@@
$\mathbb{E}V^{t+1}$ $=$ $\mathbb{E}V^{t}+\frac{\gamma^{2}}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-\frac{2\gamma}{G}\sum\limits_{i\in\mathcal{G}}\left\langle x_{i}^{t}-\overline{x}^{t},\nabla f_{i}(x_{i}^{t})\right\rangle$ (7)
$+2\gamma\left\langle\overline{x}^{t}-\overline{x}^{t},\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle$
@@
$\overset{(i)}{=}$ $\mathbb{E}\left\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right\|^{2}+\mathbb{E}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+2\mathbb{E}\left\langle g_{i}^{t}-\nabla f_{i}(x_{i}^{t}),\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\rangle$
$+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$\overset{(ii)}{=}$ $\mathbb{E}\left\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right\|^{2}+\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
$=\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\mathbb{E}\left\|g_{i}^{t}-\nabla f_{i}(x_{i}^{t})\right\|^{2}+\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}$
$\quad+\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}-2\mathbb{E}\left\|\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}g_{i}^{t}\right\|^{2}$
@@
$\overset{(i)}{\leqslant}\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\left\|\nabla f_{i}(x_{i}^{t})-\frac{1}{G}\sum\limits_{i\in\mathcal{G}}\nabla f_{i}(x_{i}^{t})\right\|^{2}+\sigma^{2},$ (8)